OpenStudy (anonymous):

Solve for x: cos^2x+sin(2x)-sinx-1=-sin^2x

OpenStudy (anonymous):

\[\sin^2x-sinx+\sin2x+\cos^2x-1=0\] \[1-sinx+\sin2x-1=0\] \[-sinx+\sin2x=0\] \[2\sin \frac{x }{ 2 }\cos \frac{ 3x }{ 2 }=0\] \[\sin \frac{x }{ 2 }\cos \frac{ 3x }{ 2 }=0\] \[\sin \frac{x }{ 2 }=0\] x/2=npi x=2npi \[\cos \frac{ 3x }{ 2 }=0\] \[\frac{ 3 }{ 2}x=\frac{ (2m+1)\pi }{ 2 }\] \[x=\frac{ (2m+1)\pi }{ 3 }\]